# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ Abstract
Peptide therapeutics, a major class of medicines, have achieved remarkable success across diseases such as diabetes and cancer, with landmark examples such as GLP-1 receptor agonists revolutionizing the treatment of type-2 diabetes and obesity. Despite their success, designing peptides that satisfy multiple conflicting objectives, such as target binding affinity, solubility, and membrane permeability, remains a major challenge. Classical drug development and target structure-based design are ineffective for such tasks, as they fail to optimize global functional properties critical for therapeutic efficacy. Existing generative frameworks are largely limited to continuous spaces, unconditioned outputs, or single-objective guidance, making them unsuitable for discrete sequence optimization across multiple properties. To address this, we present PepTune, a multi-objective discrete diffusion model for the simultaneous generation and optimization of therapeutic peptide SMILES. Built on the Masked Discrete Language Model (MDLM) framework, PepTune ensures valid peptide structures with state-dependent masking schedules and penalty-based objectives. To guide the diffusion process, we propose a Monte Carlo Tree Search (MCTS)-based strategy that balances exploration and exploitation to iteratively refine Pareto-optimal sequences. MCTS integrates classifier-based rewards with search-tree expansion, overcoming gradient estimation challenges and data sparsity inherent to discrete spaces. Using PepTune, we generate diverse, chemically-modified peptides optimized for multiple therapeutic properties, including target binding affinity, membrane permeability, solubility, hemolysis, and non-fouling characteristics on various disease-relevant targets. In total, our results demonstrate that MCTS-guided discrete diffusion is a powerful and modular approach for multi-objective sequence design in discrete state spaces.
@@ 1 Introduction
Peptides possess unique advantages as a therapeutic modality, including their ability to bind to a diverse set of binding motifs without requiring stable binding pockets, making them ideal for targeting structurally diverse protein surfaces [1, 2]. Moreover, their relatively large size and flexible backbone enable them to disrupt protein-protein interactions (PPIs) that are central to disease processes, particularly those requiring interactions with larger surface areas [3]. These attributes have driven a surge in interest, with over 33 FDA-approved therapeutic peptides introduced since 2000 and more than 170 in clinical development for diseases ranging from diabetes to cancer [2]. Specifically, peptides are behind one of the most landmark breakthroughs in recent clinical history: GLP-1 receptor agonists like semaglutide and liraglutide, which have transformed the treatment landscape for type-2 diabetes and weight loss [4, 5]. These engineered peptides have achieved remarkable efficacy, following years of meticulous structural and functional optimization [4, 5]. Their success highlights the potential of therapeutic peptides to address complex diseases where more traditional approaches, such as small molecules, often fall short [2].
@@
Generative diffusion models are considered state-of-the-art for de novo binder design, with new models such as RFpeptides even enabling the design of high-resolution macrocyclic peptides given a target structure [12]. However, structure-based models [12, 13, 14], or those that require an intermediate structure prediction step [15], rely on stable tertiary conformations of target proteins, precluding the design of peptide binders to disordered and dynamic targets, which drive a sizable portion of diseases [16]. Generative peptide design language models relying only on the target sequence, such as PepPrCLIP [17] and PepMLM [18], have demonstrated robust success on disordered and structurally-diverse targets, but their utilization of only the 20 wild-type amino acids limits these models from sampling from the space of chemically-modified or cyclic peptides, precluding exploitation of the full therapeutic potential of synthetic peptides.
Recently, discrete generative models have shown incredible promise in areas such as text generation [19, 20, 21, 22, 23, 24, 25], image synthesis [19, 21], executable code generation [23], DNA sequence optimization [26, 27, 25], and even protein design [28, 27, 24, 29], but they still face significant limitations in multi-objective-guided generation and optimization. In our case, the challenge lies in simultaneously optimizing for conflicting therapeutic properties, a critical requirement for generating clinically viable peptides [2, 30, 31]. Classifier-based and classifier-free guidance strategies have been explored to steer discrete diffusion objectives toward specific properties [29, 24, 26, 27], yet these approaches often struggle with gradient estimation and the sparsity of quality data in discrete spaces. This motivates the development of methods that can balance diverse objectives while addressing the unique challenges of discrete peptide generation.
In this work, we introduce PepTune, a multi-objective-guided discrete diffusion model for de novo peptide design. PepTune incorporates a Monte Carlo Tree Search (MCTS)-based guidance framework to condition a Masked Discrete Language Model (MDLM) [22], pre-trained on a large dataset of chemically-modified and cyclic peptides represented as Simplified Molecular Input Line Entry System (SMILES) strings [32]. However, because the vast majority of SMILES strings are neither chemically valid or nor represent synthesizable peptides, de novo sampling from the space of therapeutically relevant peptide SMILES is virtually impossible. By conditioning on synthesizability, PepTune specifically enables the generation of valid peptide SMILES containing both non-natural amino acids (nAAs) and cyclic structures while achieving Pareto-optimal solutions across multiple therapeutic properties. Our results highlight PepTune’s unique capability to balance diverse objectives, setting a new standard for property-conditioned sequence generation.
@@
    Masked Diffusion Language Model for Discrete Sampling of Peptide SMILES. We introduce the first Masked Diffusion Language Model (MDLM) with a RoFormer backbone for de novo generation of peptide SMILES sequences that is capable of generating valid and diverse chemically-modified and cyclic peptides that cannot be encoded with canonical amino acid sequences.
@@
    State-Dependent Masking Schedule. Since the presence of peptide bonds is a fundamental component of all valid peptides, we introduce a novel state-dependent masking schedule that masks peptide bond tokens at a slower rate at earlier time steps in the forward diffusion process than non-peptide bonds. We demonstrate that our masking schedule increases the weight of the negative log-loss associated with peptide bond tokens in a time-dependent manner and results in earlier unmasking of peptide bond tokens in the reverse diffusion process.
@@
    Global Sequence Invalidity Loss. Since the vast majority of SMILES sequences do not translate into valid peptides, we introduce an invalid loss that penalizes predicted token probabilities where the argmax token sequence is not a valid peptide SMILES. Since the argmax function is not differentiable, we scale the penalty score by the softmax probability of the sampled token. This objective smoothly propagates the penalty from a single discrete sequence to the smooth token probability distributions through gradient updates without the need for complex techniques like straight-through gradient estimators.
@@
    MCTS for Robust Classifier-Based Multi-Objective Guidance. We propose a novel classifier-based multi-objective guidance strategy for discrete diffusion based on the Monte-Carlo Tree Search (MCTS) algorithm that selects a sequence of optimal unmasking steps based on preceding iterations, samples a new set of unmasked sequences to compute their objective scores, and finally back-propagates the scores in the path and updates a set of global Pareto-optimal sequences. This strategy balances exploration of the peptide space through the pre-trained MDLM and multi-objective guidance through the MCTS guidance to enable the discovery of a set of valid Pareto-optimal peptide sequences. We also provide a case study for a time-dependent multi-objective guidance strategy to enable prioritization of properties.
@@
    Peptide SMILES Property Prediction Toolkit. To supplement our multi-objective guidance strategy, we train a suite of classification and regression models for property-prediction given peptide SMILES input. We train regression models to predict binding affinity to a protein target using cross-multi-head attention and membrane permeability ($\log P_{\text{exp}}$) using boosted trees. In addition, we train boosted trees for logistic regression of solubility, hemolysis, and non-fouling.
@@ 2.1 Unconditional Generator of Peptide SMILES with Masked Discrete Diffusion
Given the immense success of masked language models (MLMs) in learning bidirectional relationships of sequential data [33], including peptide-protein interaction [18], we explore the recent Masked Diffusion Language Model (MDLM) formulation (Figure 1) [22]. MDLM is a discrete diffusion architecture that leverages the MLM objective to effectively learn the clean distribution of sequences $p(\mathbf{x})$ by reconstructing clean sequences from sequences corrupted with [MASK] tokens.
@@
We specifically select the MDLM framework as our generative model to explicitly enforce SUBS parameterization such that once a token is unmasked at time $t$, it remains unmasked at the same state for the remainder of the unmasking process for $t\to 0$. Since slight modifications to the peptide sequence can result in significant alterations to its properties [34], SUBS parametrization allows us to backpropagate rewards from unmasked sequences to earlier unmasking steps that accurately reflect the favorability of the particular unmasking step in generating optimal peptide SMILES sequences.
@@
With these enhancements, the loss converged after only two epochs of training on the dynamically-batched set of 11 million peptide SMILES (Methods 4.1), with a per-token training loss of 0.832 and a validation loss of 0.880, but we trained for a total of 8 epochs (Table 1). Our optimized unconditional MDLM trained on 11 million peptide SMILES achieved a high valid peptide generation rate of 45% with a token length of 100 ($\sim$15 amino acids) and 36% with a token length of 200 ($\sim$30 amino acids) when passed through our peptide SMILES validity filter that evaluates the presence of peptide bonds and canonical and non-canonical side-chains (Methods 4.5).
@@ 2.2 PepMDLM Generates Diverse Chemically-Modified and Cyclic Peptides
To evaluate the unconditional generation quality as a benchmark for our guided generation model, we leverage the Moses metrics [36], including validity, uniqueness, diversity, and similarity to nearest neighbor (SNN) (Methods 4.5), to compare our model with an autoregressive generator of macrocyclic peptides, HELM-GPT [37]. In contrast to HELM-GPT, we use our peptide SMILES validity filter that checks not only that the SMILES string is a valid molecule but also that contains peptide bonds and valid canonical and non-canonical side-chains. Overall, our model shows increased uniqueness and diversity while lower SNN, showing PepMDLM’s capability of generating diverse peptides. Despite our lower validity, our model is unique to peptide SMILES, which has much higher granularity than HELM notation (Table 2).
@@
We show that PepMDLM generates nAAs from a collection of over 200 nAAs from SwissSidechain [40] by passing them through our peptide sequence identification function, SMILES2PEPTIDE (Algorithm 7). This demonstrates PepMDLM’s unique ability to design de novo peptides with cyclic and nAA modifications, expanding the search space of therapeutic peptides well beyond any generative model trained on canonical amino acid representations.
@@ 2.3 Multi-Objective Guidance on Therapeutic Properties for Discrete Diffusion
To generate peptides with high clinical potential, they must achieve high binding affinity with a protein target while optimizing for an array of therapeutic properties like membrane permeability to reach intracellular targets, solubility to improve drug-loading, and non-fouling and non-hemolysis to mitigate negative side-effects.
Guiding the discrete diffusion objective is challenging due to the lack of data on property-specific peptides to train large generative models and the lack of gradients of the discrete sequence space. Although existing strategies have been explored such as gradient estimation and converting to a continuous latent space [23, 26, 27], there has yet to be a robust strategy that operates directly in the discrete space and can scale for various distinct properties without sacrificing performance in any one property.
To address this gap, we introduce a classifier-based multi-objective guidance strategy to generate a set of non-dominated peptide SMILES $\mathcal{P}^{*}$ using Monte Carlo Tree Search (MCTS) (Figure 3). A sequence is non-dominated when no other sequence has a strictly larger score in one or more objectives while maintaining equal scores for the remaining objectives.
Figure 3: PepTune. PepTune is a multi-objective discrete diffusion model guided with Monte-Carlo Tree Search (MCTS). The full algorithm is detailed in Algorithm 2.
We leverage the unique capability of our unconditional MDLM model to sample from the diverse space of peptide SMILES and our trained property classifiers to search the constrained space of peptides with therapeutically optimal properties through an iterative selection, expansion, rollout, and backpropagation loop (Methods 4.4). The PepTune multi-objective guidance framework is defined as follows.
We start each iteration from a fully masked sequence $\mathbf{z}_{T}$, defining the root node of the MCTS tree. During the selection step, we traverse an optimal path through the MCTS tree defined as a series of partial unmasking steps that have been previously traversed based on a cumulative reward vector $\mathbf{U}(\mathbf{z}_{s})$ (defined in Equation 21), indicating the likelihood that the given step generates a Pareto non-dominated sequence.
Upon reaching a terminal leaf node $\mathbf{z}_{t}$ defined as a partially masked sequence at time $t>0$ that has yet to be further unmasked, we expand the leaf node to explore $M$ different possible unmasking steps by batched Gumbel unmasking (Equation 23), which applies independently sampled Gumbel noise vectors to the token probabilities predicted by the MDLM backbone for sampling each child node $\mathbf{z}_{s,i}$ for $i=1\dots M$, balancing diversity across unmasking schemes while following the predicted token distribution.
@@
After computing the reward vector for all valid sequences generated during rollout, we take the sum of the reward vectors subtracted by a penalty score proportional to the fraction of sampled SMILES that are not peptides and add the resulting $K$-dimensional vector to the cumulative reward vectors $\mathbf{W}(\mathbf{z}_{t})$ of all successor nodes leading to the root node $\mathbf{z}_{T}$, which determines the selection steps of proceeding iterations.
Even though the selection process favors high-reward unmasking steps, we show that the resulting pool of generated peptides retains similar uniqueness and diversity scores to the unconditional MDLM and the training dataset distribution (Table 3). In addition, after only 20 iterations of the MCTS search algorithm, the fraction of valid peptides reaches 100%.

### v2

@@ Abstract
Peptide therapeutics, a major class of medicines, have achieved remarkable success across diseases such as diabetes and cancer, with landmark examples such as GLP-1 receptor agonists revolutionizing the treatment of type-2 diabetes and obesity. Despite their success, designing peptides that satisfy multiple conflicting objectives, such as target binding affinity, solubility, and membrane permeability, remains a major challenge. Classical drug development and target structure-based design are ineffective for such tasks, as they fail to optimize global functional properties critical for therapeutic efficacy. Existing generative frameworks are largely limited to continuous spaces, unconditioned outputs, or single-objective guidance, making them unsuitable for discrete sequence optimization across multiple properties. To address this, we present PepTune, a multi-objective discrete diffusion model for the simultaneous generation and optimization of therapeutic peptide SMILES. Built on the Masked Discrete Language Model (MDLM) framework, PepTune ensures valid peptide structures with state-dependent masking schedules and penalty-based objectives. To guide the diffusion process, we propose a Monte Carlo Tree Search (MCTS)-based strategy that balances exploration and exploitation to iteratively refine Pareto-optimal sequences. MCTS integrates classifier-based rewards with search-tree expansion, overcoming gradient estimation challenges and data sparsity. Using PepTune, we generate diverse, chemically modified peptides optimized for multiple therapeutic properties, including target binding affinity, membrane permeability, solubility, hemolysis, and non-fouling for various disease-relevant targets. In total, our results demonstrate that MCTS-guided masked discrete diffusion is a powerful and modular approach for multi-objective sequence design in discrete state spaces.
@@ 1 Introduction
Peptides possess unique advantages as a therapeutic modality, including their ability to bind to a diverse set of binding motifs without requiring stable binding pockets, making them ideal for targeting structurally diverse protein surfaces [1, 2]. Their relatively large size and flexible backbone enable them to disrupt protein-protein interactions (PPIs) that are central to disease processes, particularly those requiring interactions with larger surface areas [3]. These attributes have driven a surge in interest, with over 33 FDA-approved therapeutic peptides introduced since 2000 and more than 170 in clinical development for diseases ranging from diabetes to cancer [2]. Specifically, peptides are behind one of the most landmark breakthroughs in recent clinical history: GLP-1 receptor agonists like semaglutide and liraglutide, which have transformed the treatment landscape for type-2 diabetes and weight loss [4, 5]. These engineered peptides have achieved remarkable efficacy, following years of meticulous structural and functional optimization [4, 5]. Their success highlights the potential of therapeutic peptides to address complex diseases where more traditional approaches, such as small molecules, often fall short [2].
@@
Generative diffusion models are considered state-of-the-art for de novo binder design, with new models such as RFpeptides even enabling the design of high-resolution macrocyclic peptides given a target structure [12]. However, structure-based models [12, 13, 14], or those that require an intermediate structure prediction step [15], rely on stable tertiary conformations of target proteins, precluding the design of peptide binders to disordered and dynamic targets, which drive a sizable portion of diseases [16]. Generative peptide design language models relying only on the target sequence, such as PepPrCLIP [17] and PepMLM [18], have demonstrated robust success on disordered and structurally diverse targets, but their utilization of only the 20 wild-type amino acids limits these models from sampling from the space of chemically-modified or cyclic peptides, precluding exploitation of the full therapeutic potential of synthetic peptides.
Recently, discrete generative models have shown incredible promise in areas such as text generation [19, 20, 21, 22, 23, 24, 25], image synthesis [19, 21], executable code generation [23], DNA sequence optimization [26, 27, 25], and even protein design [28, 27, 24, 29], but they still face significant limitations in multi-objective-guided generation and optimization. In our case, the challenge lies in simultaneously optimizing for conflicting therapeutic properties, a critical requirement for generating clinically viable peptides [2, 30, 31]. Classifier-based and classifier-free guidance strategies have been explored to steer discrete diffusion objectives toward specific properties [29, 24, 26, 27], yet these approaches often struggle with conflicting objectives, gradient estimation, and the sparsity of quality data.
In this work, we introduce PepTune, a multi-objective-guided discrete diffusion model for de novo peptide design. PepTune incorporates a novel Monte Carlo Tree Search (MCTS)-based framework for multi-objective guidance of a generative masked discrete diffusion model, pre-trained on a large dataset of chemically modified and cyclic peptides represented as Simplified Molecular Input Line Entry System (SMILES) strings [32]. However, due to the granularity of SMILES representations, the vast majority of SMILES strings are neither chemically valid nor represent synthesizable peptides. By leveraging a state-dependent masking schedule and invalidity loss function, PepTune is capable of selectively sampling from the sub-space of valid peptide SMILES containing both non-natural amino acids (nAAs) and cyclic structures while maintaining a set of Pareto-optimal solutions across multiple therapeutic properties. Our results highlight PepTune’s unique capability to balance diverse objectives, setting a new standard for property-conditioned sequence generation.
@@
    Masked Diffusion Language Model for Discrete Sampling of Peptide SMILES. We introduce the first Masked Diffusion Language Model (MDLM) [22] with a RoFormer backbone for de novo generation of peptide SMILES sequences that is capable of generating valid and diverse chemically modified and cyclic peptides that cannot be encoded with canonical amino acid sequences.
@@
    State-Dependent Masking Schedule. Since the presence of peptide bonds is a fundamental component of all valid peptides, we derive the reverse posterior and NELBO for a state-dependent masking schedule that masks peptide bond tokens at a slower rate at earlier time steps in the forward diffusion process than non-peptide bonds. This strategy increases the weight of the NELBO loss associated with peptide bond tokens in a time-dependent manner and results in earlier unmasking of peptide bond tokens in the reverse diffusion process.
@@
    Global Sequence Invalidity Loss. Since the vast majority of SMILES sequences do not translate into valid peptides, we introduce an invalid loss that penalizes predicted token probabilities where the argmax token sequence is not a valid peptide SMILES. We scale the penalty score by the softmax probability of the sampled token. This objective propagates the penalty from a single discrete sequence to the smooth token probability distributions, shifting the model parameters away from invalid predictions.
@@
    MCTS for Robust Classifier-Based Multi-Objective Guidance. We propose a classifier-based multi-objective guidance strategy for discrete diffusion based on the Monte-Carlo Tree Search (MCTS) algorithm that selects a sequence of optimal unmasking steps based on preceding iterations, samples a new set of unmasked sequences to compute their objective scores, and finally back-propagates the scores in the path and updates a set of global Pareto-optimal sequences. This strategy balances exploration of the peptide space through the pre-trained MDLM and multi-objective guidance through the MCTS guidance to enable the discovery of a set of valid Pareto-optimal peptide sequences. We also provide a case study for a time-dependent multi-objective guidance strategy to enable prioritization of properties.
@@
    Peptide SMILES Property Prediction Toolkit. To supplement our multi-objective guidance strategy, we train a suite of classification and regression models for property-prediction given peptide SMILES input. We train regression models to predict binding affinity to a protein target using cross-multi-head attention and cell membrane permeability using boosted trees. In addition, we train boosted trees for logistic regression of solubility, hemolysis, and non-fouling.
@@ 2.1 Unconditional Generator of Peptide SMILES with Masked Discrete Diffusion
Given the immense success of masked language models (MLMs) in learning bidirectional relationships of sequential data [33], including peptide-protein interaction [18], we based our unconditional generator on the Masked Diffusion Language Model (MDLM) framework (Figure 1) [22]. MDLM is a discrete diffusion architecture that leverages the MLM objective to effectively learn the clean distribution of sequences $p(\mathbf{x})$ by reconstructing clean sequences from sequences corrupted with [MASK] tokens.
@@
We enforce SUBS parametrization [22] such that once a token is unmasked at time $t$, it remains unmasked at the same state for the remainder of the unmasking process for $t\to 0$. Since slight modifications to the peptide sequence can result in significant alterations to its properties [34], SUBS parametrization allows us to backpropagate rewards from unmasked sequences to earlier unmasking steps that accurately reflect the favorability of the particular unmasking step in generating optimal peptide SMILES sequences.
@@
With these enhancements, the loss converged after only two epochs of training on the dynamically-batched set of 11 million peptide SMILES (Methods 4.1), with a per-token training loss of 0.832 and a validation loss of 0.880, but we trained for a total of 8 epochs (Table 1). Our optimized unconditional MDLM (PepMDLM) trained on 11 million peptide SMILES achieved a high valid peptide generation rate of 45% with a token length of 100 ($\sim$15 amino acids) and 36% with a token length of 200 ($\sim$30 amino acids) when passed through our peptide SMILES validity filter that evaluates the presence of peptide bonds and canonical and non-canonical side-chains (Methods 4.5).
@@ 2.2 PepMDLM Generates Diverse Chemically-Modified and Cyclic Peptides
To evaluate the unconditional generation quality as a benchmark for our guided generation model, we leverage the Moses metrics [36], including validity, uniqueness, diversity, and similarity to nearest neighbor (SNN) (Methods 4.5), to compare our model with an autoregressive generator of macrocyclic peptides, HELM-GPT [37]. In contrast to HELM-GPT, we use our peptide SMILES validity filter that checks that the SMILES string is a valid molecule and contains peptide bonds and valid natural or non-natural side chains. Overall, PepMDLM shows increased uniqueness and diversity with lower SNN, demonstrating our capability to comprehensively search the sub-space of valid peptide SMILES. Despite our lower validity, our model is unique to peptide SMILES, which has much higher granularity than HELM notation (Table 2).
@@
Similarly, PepMDLM generates valid peptides with an average slightly larger than both datasets, demonstrating our unique ability to design de novo peptides with cyclic and nAA modifications and expanding the search space of therapeutic peptides well beyond any generative model trained on canonical amino acid representations.
@@ 2.3 Multi-Objective Guidance on Therapeutic Properties for Discrete Diffusion
To generate peptides with high clinical potential, they must achieve high binding affinity with a protein target while optimizing for an array of therapeutic properties like membrane permeability to reach intracellular targets, solubility to improve drug-loading, and non-fouling and non-hemolysis to mitigate negative side-effects. Guiding the discrete diffusion objective is challenging due to the lack of data on property-specific peptides to train large generative models and the lack of gradients of the discrete sequence space. Although existing strategies have been explored such as gradient estimation and converting to a continuous latent space [23, 26, 27], there has yet to be a robust strategy that operates directly in the discrete space and can scale for various distinct properties without sacrificing performance in any one property.
To address this gap, we introduce a classifier-based multi-objective guidance strategy to generate a set of non-dominated peptide SMILES $\mathcal{P}^{*}$ using Monte Carlo Tree Search (MCTS) (Figure 3). A sequence is non-dominated when no other sequence has a strictly larger score in one or more objectives while maintaining equal scores for the remaining objectives. We leverage the unique capability of our unconditional MDLM model to sample from the diverse space of peptide SMILES and our trained property classifiers to search the constrained space of peptides with therapeutically optimal properties through an iterative selection, expansion, rollout, and backpropagation loop (Methods 4.4).
Figure 3: PepTune. PepTune is a multi-objective discrete diffusion model guided by Monte-Carlo Tree Search (MCTS). The full algorithm is detailed in Algorithm 3.
The PepTune multi-objective guidance framework is defined as follows. We start each iteration from a fully masked sequence $\mathbf{z}_{t(T)}$, defining the root node of the MCTS tree. During the selection step, we traverse an optimal path through the MCTS tree defined as a series of partial unmasking steps that have been previously traversed based on the selection score vector $\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})$ (defined in Equation 19), indicating the likelihood that the given step generates a Pareto non-dominated sequence.
Upon reaching a terminal leaf node $\mathbf{z}_{t}$ defined as a partially masked sequence at time $t>0$ that has yet to be further unmasked, we expand the leaf node to explore $M$ different possible unmasking steps by batched Gumbel unmasking (Equation 21), which applies independently sampled Gumbel noise vectors to the token probabilities predicted by the MDLM backbone for sampling each child node $\mathbf{z}_{s,i}$ for $i=1\dots M$, enforcing diverse unmasking schemes while remaining consistent with the predicted token distribution.
@@
After computing the reward vector for all valid sequences generated during rollout, we take the sum of the reward vectors subtracted by a penalty score proportional to the fraction of sampled SMILES that are not peptides and add the resulting $K$-dimensional vector to the cumulative reward vectors $\mathbf{W}(\mathbf{z}_{t})$ of all predecessor nodes on the path to the root node $\mathbf{z}_{t(T)}$, which determines the selection steps of proceeding iterations.
Even though the selection process favors high-reward unmasking steps, we show that the resulting pool of generated peptides retains similar uniqueness and diversity scores to the peptides generated by our unconditional MDLM and in the training dataset (Table 3). In addition, the fraction of valid peptides consistently reaches 100% after only 20 iterations of the MCTS search algorithm, demonstrating the effectiveness of backpropagating the classifier-based rewards and invalidity penalty.